OpenStudy (anonymous):

Please help me with this question! I know y = -1, but I do not know how to solve it! (Equation is in the replies.)

OpenStudy (anonymous):

\[y-1=4y-2/3\]

OpenStudy (anonymous):

Did it come up?

OpenStudy (anonymous):

\[y-1=\frac{ 4y-2 }{ 3 }\]

OpenStudy (anonymous):